Fix Version/s: None
Sprint:TSSW Sprint - Oct 12 - Oct 26
Team:Telescope and Site
Support the summit activity of rebooting the M2 and hexapod/rotator servers. The action of M2 server will be at 10/14 (SUMMIT-4314) and the action of hexpod/rotator server will be at 10/15 (SUMMIT-4315).
This task is in the phase 1.
We powered on the hex/rot server and Andrew saw 5 vnc servers are running now.
Andrew and I created the following tickets to ask for IT's help: IHS-4271, IHS-4272, and IHS-4274.
I can ssh in the M2 server and M2 cRIO now.
I can ssh in the hex/rot server now.
Julio will work on M2 server tomorrow. He thought he touched the BIOS with Dave in the previous time.
Based on Andrew, Julio had fixed the M2 GUI (BIOS problem). He had to disable the integrated graphics.
When we tried to execute the M2 control system in LabVIEW 2016, the application freezes and there is no response. Try to use the updated M2 control system under my account (ttsai). I tested the SAL by cpp and there should be no problem.
Andrew can run the M2 control system now. People will test the M2 mirror tomorrow.
M2 can be put into the closed-loop already. Franco and Roberto will do the following check. Andrew will focus on the rot/hex/ccw now. He and I created a diagram to show IT the network we need. IT should support this in tomorrow.
The ThinkPad and hex/rot server can ping each other already. Andrew powered on the rotator PXI controller by netbooter and I could ping the rotator PXI controller from the management PC. Based on the Andrew's feedback, it looks like there is the problem of tcp/ip connection of port 5580 (5581 looks good). The 5580 is the telemetry and 5581 is the command.
Summary of today's test:
1. We cleaned the tigervnc temporary files (X5 - X9) and made sure 5 vncserver services are runnung in the management PC.
2. We made sure only the rotator 2 and camera hexapod can be powered on. The M2 hexapod controller is in the base office.
3. We used the netstat in the management PC to make sure the connections of 5580 (tmosms0), 5581 (tmosms1), 5590, and 5591 are established.
4. We made sure they are the same in the controller.
5. We used the tcpdump command in the management PC to make sure we had the packet from controller (5580 and 5590 are the telemetry ports from the low-level controler to the management PC).
6. If we use the ThinkPad rotator/hexapod GUI, we have the error from telemetry ports.
Te-Wei and I did work on the network over this past week. We made good progress, still have a few issues to resolve. I will be creating a ticket for this soon.
Andrew powered on the M2 server and logged in as the aura user. However, we could not let the GNONE desktop to work. We tried the followings already:
We suspected that this comes from the driver. IT changed the output to VGA before in the configuration file. Andrew will create a JIRA ticket to let the IT to help. The ticket is IHS-4270.
Andrew confirmed the M2 server had connected to the control network already. However, the control network is not working yet. It might be in tomorrow.